OpenStudy (anonymous):

According to NASA, Mars is 225,000,000 km from the Sun at its closest point.

Parth (parthkohli):

Scientific notation?

Parth (parthkohli):

Shift the decimal like this: \( \color{Black}{\Rightarrow 225 \times 1,000,000 }\) \( \color{Black}{\Rightarrow 2.25 \times 100,000,000}\) \( \color{Black}{\Rightarrow 2.25 \times 10^{8} }\)

OpenStudy (aravindg):

was that what u were looking for?

OpenStudy (anonymous):

kind of in a way, yes. i remember reading that in my book, but i've not been able to get into my book this weekend